find the difference between an angel such that its supplement and two times its complement is 35°
Answer by Boreal(15235)

angle = x
complement is 90-x
supplement is 180-x
180-x-35=2(90-x)
145-x=180-2x
x=35 degrees
complement is 55
supplement is 145 degrees.
I found the difference between the supplement and twice the complement is 35 degrees. I think that is what was asked for.